# Unlocking Anticipatory Text Generation: A Constrained Approach for Faithful Decoding with Large Language Models

### v1

@@ 1 Introduction
Large language models (LLMs) exhibit impressive textual understanding and reasoning capabilities as evidenced by various studies (Brown et al., 2020; Kojima et al., 2022; OpenAI, 2022; 2023). Through the process of instruction tuning, where large models are fine-tuned on data comprising diverse tasks with specific instructions, their performance can be notably improved, even for unseen tasks. However, despite their strong abilities in text understanding and generation, undesirable behaviors such as toxicity (Hartvigsen et al., 2022) and hallucination (Ji et al., 2023) still persist. In particular, ensuring that the models’ outputs closely align with provided prompts remains a challenge. Figure 1 provides an illustration of how model-generated texts can deviate significantly from the instructions provided in their prompts, but still remain fluent and relevant.
@@
Specifically, in order to mitigate undesirable behaviors and ensure faithfulness to instructions, we propose a novel approach for text generation, by formalizing it as a problem constrained by future language generation. A future-constrained satisfaction score is incorporated for guiding the next token generation. This approach serves to steer the generation process close to desired behaviors and follow with the specified instructions. As shown in Figure 1, the future constrain score is used to choose a better next token to complete a sentence.
@@
With the future constraint satisfaction, we can search the best sequence over the infinite output space. In order to speed up the process, we present a beam-based algorithm meticulously crafted to recursively generate sequences from left to right, remarkably enhancing the efficiency and efficacy of the generation process. The experimental results exhibit desired behaviour improvements in three different tasks: keyword-constrained generation, toxicity reduction, and factual correctness in question answering. It sheds light on the pathway for achieving faithful decoding with large language models through our approach.
@@ 2 Method
Here $p(y_{t}\mid{\bm{y}}_{<t},{\bm{x}})$ represents the distribution of the next token at position $t$ given the prompt/prefix ${\bm{x}}$, and the partial output ${\bm{y}}_{<t}$. All sequential tokens are iteratively generated based on this conditional probability distribution.
@@
$\log p({\bm{y}}\mid{\bm{x}})=$ $\sum_{t}\log p(y_{t}\mid{\bm{y}}_{<t},{\bm{x}})\propto\sum_{t}\log\Bigl{(}p(y_{t}\mid{\bm{y}}_{<t})*p({\bm{x}}\mid{\bm{y}}_{<=t})\Bigr{)}$
$\approx$ $\sum_{t}\log\Bigl{(}p(y_{t}\mid{\bm{y}}_{<t},{\bm{x}})*p(C({\bm{x}})\mid{\bm{y}}_{<=t})\Bigr{)}\quad\quad\quad C({\bm{x}})\ \texttt{can be}\ {\bm{x}}$
@@ 2.1 Estimation of Future Constraint Satisfaction
In our method, we utilize future constraint satisfaction to provide guidance for text generation while ensuring the decoding efficiency of large language models. In this subsection, we introduce how to estimate the future constraint satisfaction using LLMs.
@@
$R({\bm{y}}_{<=t},C({\bm{x}}))=\frac{\log p(C({\bm{x}})\mid{\bm{y}}_{<=t},\mathrm{\textlangle SEP\textrangle})}{|C({\bm{x}})|}$ (1)
where $\mathrm{\textlangle SEP\textrangle}$ is the special token delimiting the two sequences.
@@
Note that, we rely solely on the likelihood of pretrained language models to estimate the satisfaction in this study. While this offers considerable versatility and flexibility, it might not always yield precise estimations. One can leverage fine-tuning and parameter-efficient techniques like LoRA (Hu et al., 2022) to effectively tailor the estimation process, providing more accurate and flexible assessments of constraint satisfaction. We leave this to future work.
@@ 3.1 Keyword-constrained Generation
In our initial task, we focus on lexical-constrained text generation using the CommonGen dataset (Lin et al., 2020). This task involves generating a sentence containing specific given keywords. For instance, given a set of concepts (e.g., car, drive, snow), the objective is to generate a fluent sentence that incorporates these concepts (e.g., ”I drive my car during the winter through the snow”). We evaluate the generated outputs using automatic metrics of fluency (BLEU, CIDER, etc.) and a constraint coverage score. The coverage score is calculated as the average percentage of the provided concepts present in the generated outputs.
@@
In order to check the estimation quality of future constraint satisfaction using LLMs, we create a ranking benchmark, where each sample consists of a sentence pair $({\bm{a}},{\bm{b}})$, with ${\bm{a}}$ being the sentence with a constraint $C$ and ${\bm{b}}$ without. Each ${\bm{a}}$ is derived from the development set of CommonGen, while ${\bm{b}}$ is a complete sentence generated by ChatGPT given a few prefix words from ${\bm{a}}$. We hypothesize that if this completed sentence ${\bm{b}}$ does not include all the specified concepts, it should be treated as a negative sample compared to ${\bm{a}}$.
We also investigate a distinct scenario involving a sequence pair $(\hat{{\bm{a}}},\hat{{\bm{b}}})$, where both sequences have similar lengths and are incomplete. The sole distinction between them lies in the last word, while they share the same prefix. $\hat{{\bm{a}}}$ and $\hat{{\bm{b}}}$ have the same prefix, except for the last word. Specifically, $\hat{{\bm{a}}}$ is the prefix of ${\bm{a}}$, and $\hat{{\bm{b}}}$ has the same prefix as $\hat{{\bm{a}}}$, except for the last word. The last word in $\hat{{\bm{b}}}$ is a randomly selected word from ${\bm{b}}$ [^2] Although $\hat{{\bm{a}}}$ and $\hat{{\bm{b}}}$ differ by only one word, it’s important to note that their tokenized sequences may have varying lengths. However, the difference in length is small..
@@
For each sentence pair $({\bm{a}},{\bm{b}})$, we assign a ranking accuracy score of 1 if $R({\bm{a}},C)>R({\bm{b}},C)$. Otherwise, the ranking accuracy score is 0. Figure 2 shows the ranking accuracies of keyword-constrained satisfaction estimation using various models [^3] More details about these models are in Section A.1 . High accuracies over sentence pairs are observed. However, accuracy significantly drops for prefix pairs, suggesting that satisfaction estimation for prefix pairs is considerably more challenging. Fortunately, many open LLMs still manage to achieve over 60% accuracy. Another observation is the high performance achieved by NLI-based models, despite their significantly smaller model sizes.
@@
We examine the effect of $\lambda$ in our proposed method. In Figure 3, we display the constraint coverage of sentences and BLEU-4 scores on the CommonGen development set. $\lambda=0$ corresponds to a decoding method without considering future constraint satisfaction. For $\lambda$ in the range $\lambda\in\{1,2,\dots,10\}$, our proposed method consistently achieves higher coverage scores, indicating a higher percentage of provided concepts present in the generated outputs. However, setting a very large $\lambda$ can excessively weigh on the constraint satisfaction term and hurt performance.
@@ 3.2 Toxicity Reduction
Next, we consider another task: toxicity reduction (Liu et al., 2021). Given a prompt ${\bm{x}}$, the task is to generate a fluent continuation ${\bm{y}}$ but not with a toxicity attribute. The next token is generated recursively by sampling next token probability distribution provided by LLMs. Following to the setting in  Liu et al. (2021), we use the REALTOXICITYPROMPTS benchmark (Gehman et al., 2020), generating up to 20 tokens with nucleus sampling ($p=0.9$). Following previous work (Gehman et al., 2020), we characterize generation toxicity using the toxicity score from Perspective API. Two toxicity scores are reported: 1) maximum toxicity, defined as the average maximum toxicity over 25 sampled generations, and 2) the (empirical) toxicity probability of at least 1 out of 25 generations being toxic. We also evaluate our generations for fluency, and diversity. Diversity is another metric, which is the mean number of distinct n-grams, normalized by the length of text.
In our proposed method, we reweight the top $k=50$ token logits from LLMs with our future constraint satisfaction score, then truncate the logits that are in the top-k/top-p vocabulary at each position, effectively assigning zero probability to tokens outside the vocabulary. We determine the hyperparameter $\lambda$ by evaluating its performance on a set of 50 samples.
@@
To evaluate the quality of toxicity constraint scores from LLMs, we establish our ranking benchmark. Constructing sequence pairs $({\bm{a}},{\bm{b}})$ where ${\bm{a}}$ is less toxic than ${\bm{b}}$, we utilize a file containing numerous model outputs and human-evaluated toxicity scores. [^4] The file can be accessed at https://github.com/alisawuffles/DExperts/blob/main/human_eval/toxicity/human_eval_toxicity.csv., provided by the work (Liu et al., 2021).
From the given file, we create sequence pairs $({\bm{a}},{\bm{b}})$ by employing the same prompt prefix and pairing it with two distinct annotated continuations, each having its own toxicity score. The prefix pair $(\hat{{\bm{a}}},\hat{{\bm{b}}})$ is formed using the common prefix and the first word from these two continuations.
[…]
(a) Ranking accuracy on sentence pairs $({\bm{a}},{\bm{b}})$.
(b) Ranking accuracy on prefix pairs $(\hat{{\bm{a}}},\hat{{\bm{b}}})$.
Figure 4: Accuracy of the estimation of constraint satisfaction with different models.
@@ Results.
Table 2 presents the toxicity reduction on two different LLMs (Falcon-7B-Instruct and Alpaca-7B-Instruct), which also have a minor decrease on diversity. We do not include LLaMA-2-13B-Chat because we notice that it is a low toxicity mode as shown in Touvron (2023) [^6] We also conducted tests and discovered that the average maximum toxicity score is approximately 0.135, while the average toxicity probability is close to 0.01..
@@
Toxicity (normal-$\downarrow$) Diversity (normal-$\uparrow$)
Avg. Max
Prob.
Dist-1 Dist-2 Dist-3
Falcon-7B-Instruct
Baseline 0.371
0.215
0.549 0.839 0.843
Our 0.287
0.125
0.583 0.782 0.762
Alpaca-7B-Instruct
Baseline 0.272
0.140
0.471 0.714 0.745
Our 0.235
0.108
0.471 0.584 0.574
@@ 3.3 Factual Question Answering
Hallucination is a notable issue associated with large language models, despite their ability to generate coherent and fluent output. Providing accurate answers supported by concrete evidence is crucial, and mitigating hallucination is key to achieving this goal.
We use the dateset ALCE (Gao et al., 2023) as factual question answering This benchmark provides a set of retrieved passages, denoted as $D=\{D_{1},D2,\dots\}$, for each question $q$. Additionally, the dataset offers correctness evaluation through multiple short answers in ASQA (Stelmakh et al., 2022) and three “sub-claims” for ELI5 (Fan et al., 2019).
In ASQA, correctness is determined by calculating the recall of correct short answers. This is achieved by verifying whether the short answers provided by the dataset are exact substrings of the generated response. On the other hand, for the long-form QA task ELI5, correctness is measured by the ratio of model outputs that entail the three provided ”sub-claims”.
@@
For each evidence, if the factual constraint estimation score is higher for the supported claim compared to the unsupported claim with respect to the evidence, we assign an accuracy score of 1. Otherwise, if $R({\bm{a}},\mathrm{evidence})\leq R({\bm{b}},\mathrm{evidence})$, the accuracy score is 0. Table 3.3 displays the accuracies on our constructed factual correctness ranking benchmark. We can see that several open LLMs achieve more than 60% accuracy [^7] We noticed an usual trend in the performance of the llama-1 family model. Interestingly, we found that their performance on the Fever ranking part worsened with an increase in model size..

### v3

# Unlocking Anticipatory Text Generation: A Constrained Approach for Large Language Models Decoding

@@ Abstract
Unlocking Anticipatory Text Generation: A Constrained Approach for Large Language Models Decoding
Lifu Tu,  Semih Yavuz, Jin Qu, Jiacheng Xu, Rui Meng, Caiming Xiong, Yingbo Zhou Salesforce AI Research ltu@salesforce.com
@@ 1 Introduction
Large language models (LLMs) exhibit impressive textual understanding and reasoning capabilities as evidenced by various studies (Brown et al., 2020; Kojima et al., 2022; OpenAI, 2022, 2023). Through the process of instruction tuning, where large models are fine-tuned on data comprising diverse tasks with specific instructions, their performance can be notably improved, even for unseen tasks. However, despite their strong abilities in text understanding and generation, undesirable behaviors such as toxicity (Hartvigsen et al., 2022) and hallucination (Ji et al., 2023) still persist. In particular, ensuring that the models’ outputs closely align with provided prompts remains a challenge. Figure 1 provides an illustration of how model-generated texts can deviate significantly from the instructions provided in their prompts, but still remain fluent and relevant.
@@
Specifically, in order to mitigate undesirable behaviors and ensure faithfulness to instructions, we propose a novel approach for text generation (Section 2), by formalizing it as a problem constrained by future language generation. A future-constrained satisfaction score is incorporated for guiding the next token generation. This approach serves to steer the generation process close to desired behaviors and follow with the specified instructions. As shown in Figure 1, the future constrain score is used to choose a better next token to complete a sentence.
@@
With the future constraint satisfaction, we can search the best sequence over the infinite output space. In order to speed up the process, we present a beam-based algorithm meticulously crafted to recursively generate sequences from left to right, remarkably enhancing the efficiency and efficacy of the generation process. The experimental results (Section 3) exhibit desired behaviour improvements in three different tasks: keyword-constrained generation, toxicity reduction, and factual correctness in question answering. We also conduct speed and human evaluation (Section 4) of our approach. The decoding time slowdown linear with the number of candidates at each step [^1] Future work can focus on enhancing constraint satisfaction estimation and reducing candidate numbers to boost speed and performance.. It sheds light on the pathway for achieving faithful decoding with large language models through our approach.
@@ 2 Method
Here $p(y_{t}\mid{\bm{y}}_{<t},{\bm{x}})$ represents the distribution of the next token at position $t$ given the prompt/prefix ${\bm{x}}$, and the partial output ${\bm{y}}_{<t}$. All sequential tokens are generated iteratively based on this conditional probability distribution.
There are several popular deterministic decoding methods such as greedy decoding and beam search, as well as non-deterministic sampling methods like temperature sampling, nucleus sampling (Holtzman et al., 2020), and top-k sampling. In this context, our focus primarily revolves around deterministic decoding techniques.
@@
$\log$ $p({\bm{y}}\mid{\bm{x}})=\sum_{t}\log p(y_{t}\mid{\bm{y}}_{<t},{\bm{x}})$
$\propto$ $\sum_{t}\log\Bigl{(}p(y_{t}\mid{\bm{y}}_{<t})*p({\bm{x}}\mid{\bm{y}}_{<=t})\Bigr{)}$
$\approx$ $\underbrace{\sum_{t}\log\Bigl{(}p(y_{t}\mid{\bm{y}}_{<t},{\bm{x}})*p(C({\bm{x}})\mid{\bm{y}}_{<=t})\Bigr{)}}_{C({\bm{x}})\ \texttt{can be}\ {\bm{x}}}$
@@ 2.1 Estimation of Future Constraint Satisfaction
In our method, we utilize future constraint satisfaction to provide guidance for text generation while ensuring the decoding efficiency of LLMs. In this subsection, we introduce how to estimate the future constraint satisfaction using LLMs.
@@
$R({\bm{y}}_{<=t},C({\bm{x}}))=\frac{\log p(C({\bm{x}})\mid{\bm{y}}_{<=t},\textlangle\mathrm{SEP}\textrangle)}{|C({\bm{x}})|}$ (1)
where $\mathrm{\textlangle SEP\textrangle}$ is the special token delimiting the two sequences [^2] We set it as the end of sequence token..
@@
In section 3, we exemplify how the proposed method can be applied to specific NLP problems. Note that, we use the likelihood of pretrained language models to estimate the satisfaction in this study. While this offers considerable versatility and flexibility, it might not always yield precise estimations. One can leverage fine-tuning and parameter-efficient techniques like LoRA (Hu et al., 2022) to effectively tailor the estimation process, providing more accurate and flexible assessments of constraint satisfaction. We leave this to future work.
@@ 3.1 Keyword-constrained Generation
In our initial task, we focus on lexical-constrained text generation using the CommonGen dataset (Lin et al., 2020). This task involves generating a sentence containing specific given keywords. For instance, given a set of concepts (e.g., car, drive, snow), the objective is to generate a fluent sentence that incorporates these concepts (e.g., "I drive my car during the winter through the snow"). We evaluate the generated outputs using automatic metrics of fluency (BLEU, CIDER, etc.) and a constraint coverage score. The coverage score is calculated as the average percentage of the provided concepts present in the generated outputs.
@@
In order to check the estimation quality of future lexical-constraint satisfaction using LLMs, we create a ranking benchmark, where each sample consists of a sentence pair $({\bm{a}},{\bm{b}})$, with ${\bm{a}}$ being the sentence with a constraint $C$ and ${\bm{b}}$ without. Each ${\bm{a}}$ is derived from the development set of CommonGen, while ${\bm{b}}$ is a complete sentence generated by ChatGPT given a few prefix words from ${\bm{a}}$. We hypothesize that if this completed sentence ${\bm{b}}$ does not include all the specified concepts, it should be treated as a negative sample compared to ${\bm{a}}$.
We also investigate a distinct scenario (prefix pairs) involving a sequence pair $(\hat{{\bm{a}}},\hat{{\bm{b}}})$, where both sequences have similar lengths and are incomplete. The sole distinction between them lies in the last word, while they share the same prefix. $\hat{{\bm{a}}}$ and $\hat{{\bm{b}}}$ have the same prefix, except for the last word. Specifically, $\hat{{\bm{a}}}$ is the prefix of ${\bm{a}}$, and $\hat{{\bm{b}}}$ has the same prefix as $\hat{{\bm{a}}}$, except for the last word. The last word in $\hat{{\bm{b}}}$ is a randomly selected word from ${\bm{b}}$ [^5] Although $\hat{{\bm{a}}}$ and $\hat{{\bm{b}}}$ differ by only one word, it’s important to note that their tokenized sequences may have varying lengths. However, the difference in length is small..
@@
For each sentence pair $({\bm{a}},{\bm{b}})$, we assign a ranking accuracy score of 1 if $R({\bm{a}},C)>R({\bm{b}},C)$. Otherwise, it is 0. Figure 2 shows the ranking accuracies of keyword-constrained satisfaction estimation using various models [^6] For more detailed information about these models, please refer to the Appendix in Section .1.. High accuracies over sentence pairs are observed. However, accuracy significantly drops for prefix pairs, suggesting that satisfaction estimation for prefix pairs is considerably more challenging. Fortunately, many open LLMs still manage to achieve over 60% accuracy. Another observation is high performance with NLI-based models, despite significantly smaller model sizes.
@@
In Figure 3, we display the constraint coverage and BLEU-4 scores on the CommonGen development set with different $\lambda$. $\lambda=0$ corresponds to a decoding method without considering future constraint satisfaction. For $\lambda$ in the range $\lambda\in\{1,2,\dots,10\}$, our proposed method consistently achieves higher coverage scores, indicating a higher percentage of provided concepts present in the generated outputs. However, setting a large $\lambda$ can excessively weight on the constraint satisfaction term and hurt performance.
@@ Results.
Comparison with NeuroLogic-A*.
No external modules and no training is used in our method, so greedy decoding, beam search are the chosen deterministic decoding baseline. NeuroLogic-A* (Lu et al., 2022b) is another baseline, however, it only applied into lexical-constrained generation tasks. We adopt the work of NeuroLogic-A* into LLMs decoding, have our own implementation, and report the results:Time and performance). We do the comparison on the lexical-constrained generation task. The instruction inputs are the same for different decoding methods. As shown in Figure 4, Our proposed method delivers results comparable to NeuroLogic-A*, but with significantly higher speed. Additionally, our method extends its utility beyond lexical constraints, encompassing applications such as toxicity reduction, ensuring factual correctness in question-answering tasks, and more. Further application results are detailed in the following sections.
Figure 4: Speed ( inference time per example ) and performance (Coverage score) of different decoding methods (with the same batch size 1 and beam size 5.). Falcon-7B-Instruct is used in this experiment. 1 A100 with 40G is used.
@@ 3.2 Toxicity Reduction
Next, we consider another task: toxicity reduction (Liu et al., 2021). Given a prompt ${\bm{x}}$, the task is to generate a fluent continuation ${\bm{y}}$ but not with a toxicity attribute. The next token is generated recursively by sampling next token probability distribution provided by LLMs. Following to the setting in  Liu et al. (2021), we use the REALTOXICITYPROMPTS benchmark (Gehman et al., 2020), generating up to 20 tokens with nucleus sampling ($p=0.9$). Following previous work (Gehman et al., 2020), we characterize generation toxicity using Perspective API. We report maximum toxicity, toxicity probability, and diversity [^7] More details are in the appendix .8..
@@
To evaluate the quality of toxicity constraint scores from LLMs, we establish our ranking benchmark. Constructing sequence pairs $({\bm{a}},{\bm{b}})$ where ${\bm{a}}$ is less toxic than ${\bm{b}}$, we utilize a file containing numerous model outputs and human-evaluated toxicity scores. [^8] The file can be accessed at https://github.com/alisawuffles/DExperts/blob/main/human_eval/toxicity/human_eval_toxicity.csv., provided by the work (Liu et al., 2021). From the given file, we create sequence pairs $({\bm{a}},{\bm{b}})$ by employing the same prompt prefix and pairing it with two distinct annotated continuations, each having its own toxicity score. The prefix pair $(\hat{{\bm{a}}},\hat{{\bm{b}}})$ is formed using the common prefix and the first word from these two continuations.
For a given prompt ${\bm{x}}$, the description of the toxicity constraint we used is $C({\bm{x}})=$“This will be a rude, disrespectful, or unreasonable comment.”. We assign a ranking accuracy score of 1 if $R({\bm{a}},C({\bm{x}}))>R({\bm{b}},C({\bm{x}}))$, otherwise 0. Figure 5 shows ranking accuracies [^9] We observe that certain pairs have nearly identical toxicity constraint scores, and we did not categorize them as incorrect. of various LLMs [^10] For more detailed information about these models, please refer to the Appendix in Section .1. on the toxicity ranking benchmark. Most open LLMs demonstrate an accuracy surpassing 50%, which represents the performance of random guessing. Particularly, the model Falcon-7B-Instruct exhibits superior performance. However, several models achieve an accuracy exceeding 60%, indicating potential for improvement in the future.
Figure 5: Accuracy of the estimation of constraint satisfaction with different pretrained LLMs.
@@ Results.
In our proposed method, we reweight the top $k=50$ token logits from LLMs with our future constraint satisfaction score, then truncate the logits that are in the top-k/top-p vocabulary at each position, effectively assigning zero probability to tokens outside the vocabulary. We determine the hyperparameter $\lambda$ by evaluating its performance on a set of 50 randomly selected samples. Table 2 presents the toxicity reduction on two different LLMs (Falcon-7B-Instruct and Alpaca-7B-Instruct), which also have a minor decrease on diversity. We do not include LLaMA-2-13B-Chat because we notice that it is a low toxicity mode as shown in Touvron (2023) [^11] We also conducted tests and discovered that the average maximum toxicity score is approximately 0.135, while the average toxicity probability is close to 0.01..
Toxicity ($\downarrow$) Diversity ($\uparrow$)
Avg. Max Prob. Dist-1 Dist-2 Dist-3
Falcon-7B-Instruct
Baseline 0.371 0.215 0.549 0.839 0.843
Our 0.287 0.125 0.583 0.782 0.762
Alpaca-7B-Instruct
Baseline 0.272 0.140 0.471 0.714 0.745
Our 0.235 0.108 0.471 0.584 0.574
@@ 3.3 Factual Question Answering
Hallucination is a notable issue associated with large language models, despite their ability to generate coherent and fluent output. Providing accurate answers supported by concrete evidence is crucial, and mitigating hallucination is key to achieving this goal. We use the dateset ALCE (Gao et al., 2023) as factual question answering This benchmark provides a set of retrieved passages, denoted as $D=\{D_{1},D2,\dots\}$, for each question $q$. Additionally, the dataset offers correctness evaluation through multiple short answers in ASQA (Stelmakh et al., 2022) and three “sub-claims” for ELI5 (Fan et al., 2019).
In ASQA, correctness is determined by calculating the recall of correct short answers. This is achieved by verifying whether the short answers provided by the dataset are exact substrings of the generated response. On the other hand, for the long-form QA task ELI5, correctness is measured by the ratio of model outputs that entail the three provided "sub-claims".
@@
For each evidence, if the factual constraint estimation score is higher for the supported claim compared to the unsupported claim with respect to the evidence, we assign an accuracy score of 1. Otherwise, if $R({\bm{a}},\mathrm{evidence})\leq R({\bm{b}},\mathrm{evidence})$, the accuracy score is 0. Table 3.3 displays the accuracies on our constructed factual correctness ranking benchmark. We can see that several open LLMs [^12] For more detailed information about these models, please refer to the Appendix in Section .1. achieve more than 60% accuracy [^13] We noticed an usual trend in the performance of the llama-1 family model. Interestingly, we found that their performance on the Fever ranking part worsened with an increase in model size..